# Sub-shot noise sensitivity via deformed four-headed kitten states

### v2

@@ Abstract
We explore nonclassical effects in the phase space of a four-headed kitten state (a superposition of two Schrödinger kitten states) induced by photon addition and subtraction operations applied in different sequences. We investigate two scenarios: in the first, photon addition is applied to the state, followed by photon subtraction, while in the second, the order of operations is reversed. We demonstrate that applying multiphoton operations to the state results in notable nearly isotropic sub-Planck structures, with the characteristics of these structures being influenced by the photon addition and subtraction. We observe that adding photons increases the average photon number, while photon subtraction reduces it in the first case but has no effect in the second. Increasing the number of added photons compresses the sub-Planck structures in both cases. Photon subtraction, however, has the opposite effect on the sub-Planck structures in the first case and no effect in the second, although it may improve their isotropy at optimal settings. The presence of the sub-Planck structures in our states leads to improved sensitivity to displacements, exceeding the standard quantum limit, as verified across all the depicted scenarios.
@@ I Introduction
Coherent states were first introduced by Schrödinger in 1926 [1, 2], and the concept was further developed in quantum optics by Glauber in 1963 [3]. Quantum superposition phenomena have been intensively investigated within the framework of the harmonic oscillator, leading to the creation of intriguing nonclassical states [4, 5]. The nonclassical nature of quantum states is revealed through nonclassical phase-space features, often visualized using the Wigner function [6, 7]. Nonclassical states are considered valuable resources for continuous variable quantum information processing, particularly for sensing and metrology applications [8, 9, 10, 11, 12]. A critical factor in determining the utility of a quantum state for such applications is the degree to which it becomes distinguishable from the initial state after a small displacement, which is influenced by the smallest phase-space feature of a quantum state and measures sensitivity to displacement (perturbation) [13, 14, 15]. Developing quantum states with finer phase-space features is essential for enhancing their capacity to measure smaller scale displacements, and quantum states with finer phase-space features are highly valuable for advancing quantum measurement techniques and improving the performance of quantum technologies [16].
The phase-space volume occupied by a coherent state adheres to the Planck scale $\hbar$ [17], while also preserving its sensitivity to displacements at this standard limit [18, 19]. It has been shown that specific chaotic quantum systems [13, 20] are able to produce phase-space features with volumes far below this standard norm. A macroscopic cat state [21, 22], which is the superposition of two distinct coherent states, is a prominent example of nonclassical states and has shown more finer phase-space features and greater sensitivity compared to a coherent state [23]. The notion of coherent-state superpositions has been evolved to a generalized form of macroscopic cat states [24, 25, 13, 11]. Compass states [26, 27, 28, 29, 30, 31, 32, 33, 34, 23, 35, 36, 37, 25, 38, 13] (a superposition of four coherent states) have been shown to exhibit improved phase-space characteristics compared to their precursors, namely coherent states and cat states. In particular, it has been found that these states hold sub-Planck scale features (dimension below than the Planck scale) and hold sensitivity to displacements greater than the standard quantum limit, which make them a potential candidate for quantum metrology applications [39]. Both theoretical [40, 41, 42, 43, 44] and experimental [45, 46, 47, 48] approaches have been employed to generate catlike states.
@@
Multiphoton operations applied to a quantum state provide innovative techniques for controlling fundamental nonclassical characteristics [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60]. For example, it has been shown that photon subtraction (or addition) from ordinary squeezed-vacuum states may develop quantum states similar to those of cat states [50, 53, 58, 52]. In this work, we consider a four-headed kitten state, which is the superposition of two kitten states deficient in nonclassical phase-space attributes. Photon addition and subtraction operations are applied on this state in different orders to build new quantum states, which exhibit intriguing phase-space characteristics similar to a compass state [13]. In our first case, we apply photon addition followed by photon subtraction to the state, while in the second case, the order of operations is reversed, with photon subtraction applied first and then photon addition.
@@
Our investigation utilizes phase-space formalism [7], incorporating Wigner function analysis and photon number distributions. We compare the states to explore the effects of multiphoton operations and the optimal selection of parameters, followed by an analysis of their sensitivity to displacements. The analysis concludes with a thorough discussion of the physical significance and implications of each scenario. Specifically, we show that multiphoton operations applied on the four-headed kitten state lead to quantum states with refined phase-space characteristics, and interestingly, our proposed cases also hold sub-Planck phase-space features whose occurence is now affilated with the number of added and subtracted photons. Furthermore, we show that our proposed instances hold sensitivity to phase-space displacements greater than the standard quantum limit, making them a promising choice for quantum sensing applications. Our observations indicate that as the number of added photons increases, finer sub-Planck structures develop in our cases. Conversely, photon subtraction in the first scenario causes the sub-Planck structures to expand. In the second scenario (subtraction followed by addition), photon subtraction becomes ineffective, but in this case, an improvement in the isotropy (directional invariance in phase space) of the sub-Planck structures is observed for certain photon subtraction choices. These findings align with the behavior observed in the sensitivity of our quantum states.
Our paper is structured as follows: §II reviews the basics of sub-Planck structures and their impact on sensitivity to displacement, illustrated with the Zurek compass state. §III discusses deformed coherent states and our proposed quantum states, offering a detailed analysis to identify the corresponding phase-space characteristics. §IV presents phase-space sensitivity contained by our proposed quantum states. §V provides a compact summary of our findings, and §VI provides the main conclusions of our results.
Figure 1: The Wigner function of the compass state. (a) $c_{0}=1$, (b) $c_{0}=5$, and (c) $c_{0}=8$.
@@ II Concepts and perspectives
The quantum uncertainty principle for position and momentum, expressed through the commutation relation $[\hat{x},\hat{p}]:=\text{i}\hbar$ with $\hat{x}$ and $\hat{p}$ being the position and momentum operators, respectively, imposes constraints on the dimension of a phase-space structure. Specifically, it dictates that the product of the uncertainties in position ($\Delta x$) and momentum ($\Delta p$) satisfies $\Delta x\Delta p\geq\nicefrac{{\hbar}}{{2}}$ [18, 19]. This implies that the dimension of a phase-space feature is constrained by this standard limit; if a phase-space feature has dimensions below this threshold, it is considered unphysical. This general assumption is challenged by Zurek [13], who demonstrates that the sub-Planck structures in the compass state significantly influence the phase-space properties of the states. These spotty features have been found crucial in enhancing sensitivity to displacements, and this improvement in sensitivity is directly dedicated to the sub-Planck scale structures in the state. In this section, we primarily focus on reviewing the key concept of sub-Planck structures and their critical role in enhancing sensitivity to displacements.
@@ II.1 Phase space and sensitivity
A Schrödinger coherent state can be expressed as a displaced vacuum state |$\ket{\alpha}:=\hat{D}(\alpha)\ket{0}$ with $\hat{D}(\alpha):=\exp(\alpha\hat{a}^{\dagger}-\alpha^{*}\hat{a})$ the displacement operator [61], where $\alpha\in\mathbb{C}$, and $\hat{a}$ ($\hat{a}^{\dagger}$) are annihilation (creation) operators. Coherent states are the most classical pure states of light [17], but their superposition may exhibit nonclassical aspects due to quantum interference [62]. The Wigner function denoted by $W_{\hat{\rho}}(\bm{\beta})$ with $\bm{\beta}:=(x,p)^{\top}$ constitutes phase space of a quantum state $\hat{\rho}$ [7], where $x$ and $p$ are the position and momentum pairs, respectively.
The Wigner function is one of the primary tools for analyzing the nonclassical characteristics of a quantum state [7]. Mathematically, the Wigner function can also be expressed as [63]
@@
The Schrödinger cat state is the superposition of two distinguishable coherent states, and one of the simplest examples of such states is the even cat state [64], which is denoted as
|$\ket{\psi}:=\frac{1}{(2+2\text{e}^{-2|\alpha|^{2}})^{\nicefrac{{1}}{{2}}}}\left[\ket{\alpha}+\ket{-\alpha}\right].$ (2)
A cat state is visible when the parameter $\alpha$ is high enough to maintain the orthogonality criterion $\braket{\alpha}{-\alpha}\approx 0$. If the constituent coherent states in this superposition are not distinguishable, the resulting states is known as Schrödinger-kitten states [52].
We now discuss the concept of sensitivity to phase-space displacement. For a pure quantum state |$\ket{\psi}$, this sensitivity can be mathematically determined by evaluating the overlap function between a quantum state and its slightly displaced version. This involves calculating how much one state resembles another when displaced in phase space, which provides insight into how precisely the quantum state can detect or respond to changes in its phase-space configuration. Mathematically, this sensitivity can be determined by using [14]
@@
with |$\ket{\psi^{\prime}}:=\hat{D}(\delta)\ket{\psi}$. If $S_{\ket{\psi}}(\delta)=0$, then a state and its displaced counterpart are orthogonal for the displacement $\delta$. The overlap $S_{\ket{\psi}}(\delta)$ with $\delta:=(\delta x,\delta p)^{\top}$, where $\delta x$ and $\delta p$ are values of the displacements applied along $x$ and $p$ directions in the phase space, respectively. The infinitesimal perturbation $\delta$, which makes the perturbed state quasi-orthogonal with the initial, provides information on the sensitivity to displacements. Smaller values of $\delta$ indicate greater sensitivity to displacements. This concept can be understood by considering a scenario where a signal intended for detection is linearly coupled to a harmonic oscillator. The oscillator measures a displacement that is proportional to the strength of the signal. In this context, the sensitivity of the oscillator to these displacements directly impacts its capacity to resolve the signal. Thus, quantum states with higher sensitivity are capable of detecting weaker signals with greater precision [26].
The Wigner function of a macroscopic cat state typically manifests as two distinct Gaussian peaks in phase space with an oscillatory interference pattern, where each peak corresponds to a coherent state; this quantum state is simply called a macroscopic cat state [65, 66]. In contrast, for a compass state, the Wigner function now holds four Gaussian peaks, a catlike interference pattern, and central sub-Planck structures [13], may also be classified as a four-component cat state [66, 67], aligning with the concept of multichotomous cat states [24]. Cat states, such as the one presented in Eq. (2), do not exhibit sub-Planck features as their interference phase-space features are not limited in all directions of phase space [23].
Figure 2: Central interference of the compass state. (a) $c_{0}=5$ and (b) $c_{0}=8$.
II.2 Sub-Planckian sensitivity
@@
which exhibits Gaussian form; hence, coherent states are types of Gaussian and nonclassical quantum states [17]. Moreover, the phase-space structure of a coherent state follows the minimal limit set by the uncertainty principle, often referred to as the Planck action in phase space. This implies that phase-space size of a coherent sets the minimal norm, and the sub-Planck structure is below this limit and can be limited as much as desired by varying the controlling parameter, whereas all of these characteristics are missing in the coherent [23].
@@
The sensitivity of a coherent state to displacements in phase space is described by the function $S_{\ket{\alpha}}(\delta):=\text{e}^{-|\delta|^{2}}$. This overlap tends to zero for the displacement $|\delta|>1$, implying that the sensitivity of a coherent state falls at the standard limits. The phase-space structure of a coherent state and its sensitivity adhere precisely to the standard quantum mechanical limits. This means that the coherent state achieves the theoretical minimum uncertainty allowed by the Heisenberg uncertainty principle, reflecting the optimal balance between precision in position and momentum measurements. Consequently, in our analysis, we evaluate each example by comparing it against these established norms. This involves assessing how each example measures up to the theoretical benchmarks and standard limits, allowing us to understand their relative performance and behavior in relation to these reference points.

### v3

# Sub-shot-noise sensitivity via superpositions of two deformed kitten states

@@ Abstract
In the present work we explore nonclassical effects in the phase space of two superposed kitten states induced by photon addition and subtraction operations applied in different sequences. We investigate two scenarios: In the first, photon addition is applied to the state, followed by photon subtraction, while in the second, the order of operations is reversed. We demonstrate that applying multiphoton operations to the state results in notable nearly isotropic sub-Planck structures, with the characteristics of these structures being influenced by the photon addition and subtraction. Increasing the number of added photons compresses the sub-Planck structures in both cases. Photon subtraction, however, has the opposite effect on the sub-Planck structures in the first case and no effect in the second. Furthermore, we observe that the optimal choices of multiphoton operations lead to improved isotropy of sub-Planck structures in our cases. The presence of the sub-Planck structures in our states leads to improved sensitivity to displacements, exceeding the standard quantum limit, as verified across all the depicted scenarios.
@@ I Introduction
Coherent states were first introduced by Schrödinger [1, 2], and the concept was further developed in quantum optics by Glauber [3]. Quantum superposition phenomena have been intensively investigated within the framework of the harmonic oscillator, leading to the creation of intriguing nonclassical states [4, 5]. The nonclassical nature of quantum states is revealed through nonclassical phase-space features, often visualized using the Wigner function [6, 7]. Nonclassical states are considered valuable resources for continuous variable quantum information processing, particularly for sensing and metrology applications [8, 9, 10, 11, 12]. A critical factor in determining the utility of a quantum state for such applications is the degree to which it becomes distinguishable from the initial state after a small displacement, which is influenced by the smallest phase-space feature of a quantum state and measures sensitivity to displacement (perturbation) [13, 14, 15].
The phase-space volume occupied by a coherent state adheres to the Planck scale $\hbar$ [16], while also preserving its sensitivity to displacements at this standard limit [17, 18]. Developing quantum states with finer phase-space features is essential for enhancing their capacity to measure smaller scale displacements, and quantum states with finer phase-space features are highly valuable for advancing quantum measurement techniques and improving the performance of quantum technologies [19], which specifically translated to sensing platforms based on superconducting circuits [20, 21, 22], trapped ion systems [23, 24], and magnon sensing [25, 26], to name a few. Nonclassical states are also crucial to detecting weak forces (tidal gravitational forces) [27, 28, 29, 30, 31]. In these contexts, minimal uncertainty states [16], such as coherent states, define the classical limit for force detection. This limit can be surpassed by quantum states exhibiting finer phase-space features [27, 32], with certain chaotic quantum systems demonstrating states whose phase-space volumes are significantly smaller than those of coherent states [13, 33].
The effectiveness of weak-force detection techniques appears to be directly impacted by the degree of phase-space feature fineness in quantum states. For example, a macroscopic cat state (superpositions of two distinct coherent states) [34, 35] exhibits finer phase-space structures than coherent states; that is, such features are constrained along a specific direction in the phase space, thus holding the capacity to detect the directional weak forces [28, 29, 27]. The notion of coherent-state superpositions has been evolved to a generalized form of macroscopic cat states [36, 37, 13, 11], and quantum compasses [28, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 37, 51, 13] are particular examples of such generic superpositions and are notable in that they exhibit improved phase-space characteristics compared to their precursors, namely coherent states and cat states. In particular, it has been found that these states contain sub-Planck scale features (dimension below the Planck scale) [13], which are extremely fragile against perturbation and exhibit sensitivity to displacements greater than the standard quantum limit, making them a potential candidate for quantum metrology applications [28, 29, 52]. Sub-Planck features have been found in entangled Schrödinger cat states created in a two-cavity setup [53], extending the notion of sub-Planck structures to higher-dimensional cat states.
Sub-Planck features, also known as blind spots [54], are constrained along all phase-space directions, allowing for the simultaneous measurement of conjugate variables (for a specific case, position, and momentum) with maximum precision [28]. These microscopic characteristics have been further extended to precisely determine the displacement of a mechanical oscillator or microwave field in any arbitrary direction within phase space [29, 28], making them superior to the cat state features. In addition to such specific cases, nonclassical states [55, 56], particularly compass states [52], have also played a crucial role as probes for enhanced phase and displacement estimation. In quantum communication channels, the robustness of the information carrier to noise is crucial [57, 58, 59, 60], and the use of finer-scale characteristics of the Wigner function has been shown to improve the effectiveness of the communication protocol, resulting in high-fidelity continuous-variable teleportation [61]. The effectiveness of finer phase-space characteristics at the sub-Planck level is crucial in continuous-variable quantum information ranging from quantum estimation to communication, and this work particularly focuses on the sub-Planck structures that are closely tied to the compasslike state of the harmonic oscillator.
@@
Both theoretical [62, 63, 64, 65, 66] and experimental [67, 68, 69, 70] approaches have been employed to generate catlike states. Multiphoton operations applied to a quantum state provide innovative techniques for controlling fundamental nonclassical characteristics [71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84]. For example, it has been shown that photon subtraction (or addition) from ordinary squeezed-vacuum states may develop quantum states similar to those of cat states [73, 76, 81, 75]. In this work, we consider a four-headed kitten state, which is the superposition of two kitten states deficient in nonclassical phase-space attributes. Photon addition and subtraction operations are applied on this multi-headed kitten state in different orders to build new quantum states, which exhibit intriguing phase-space characteristics similar to a compass state [13]. In our first case, we apply photon addition followed by photon subtraction to the state, while in the second case, the order of operations is reversed, with photon subtraction applied first and then photon addition.
@@
Contribution. Our investigation utilizes phase-space formalism [7], incorporating Wigner function analysis and photon number distributions, and we compare the original state to its deformed variants acquired from multiphoton operations by performing an overlap function between the states. This comparison investigates the ranges of parameters where the effects of multiphoton operations are highest and gives the ideal parameter selection, and we work in the region where our proposed states tend to be distinct from the original state. We also analyze displacement sensitivity for each case and present a thorough discussion of the physical significance and implications. Specifically, we show that multiphoton operations applied on the four-headed kitten state lead to quantum states with refined phase-space characteristics, and interestingly, our proposed cases also exhibit sub-Planck phase-space features whose occurrence is now affiliated with the number of added and subtracted photons. Furthermore, we show that our proposed quantum states exhibit sensitivity to phase-space displacements greater than the standard quantum limit, making them a promising choice for quantum sensing applications.
Our observations indicate that as the number of added photons increases, finer sub-Planck structures develop in the phase space of our indicated cases. In contrast, in the first case, when addition is applied prior to subtraction, corresponding photon subtraction in this case leads the sub-Planck structures to grow in size, resulting in the loss of the sub-Planck scale characteristic. In the second scenario of subtraction followed by addition, the photon subtraction operation is ineffective. In our cases, an improvement in the isotropy (directional invariance in phase space) of the sub-Planck structures is observed for certain photon subtraction choices. These findings align with the behavior observed in the sensitivity of our quantum states.
Organization. §II discusses the impact of phase-space features, particularly sub-Planck structures, on sensitivity to displacement. This analysis is carried out by comparing cat states and compass states in terms of their practical applications in estimation schemes. §III discusses the deformed version of quantum states; specifically, our focus is on the photon-varied version of coherent states and their derivatives, such as multicomponent kitten states, with a particular focus on their sub-Planck phase-space characteristics and variations over involved parameters and multiphoton operations. §IV analyzes how the sub-Planckian features in our cases have the impact on the enhancement in phase-space sensitivity to displacements. §V provides a compact summary of our findings and §VI provides the main conclusions of our results.
Figure 1: Wigner function of the compass state. for (a) $c_{0}=1$, (b) $c_{0}=5$, and (c) $c_{0}=8$.
@@ II Concepts and perspectives
The quantum uncertainty principle for position and momentum, expressed through the commutation relation $[\hat{x},\hat{p}]:=\text{i}\hbar$, with $\hat{x}$ and $\hat{p}$ the position and momentum operators, respectively, imposes constraints on the dimension of a phase-space structure. Specifically, it dictates that the product of the uncertainties in position ($\Delta x$) and momentum ($\Delta p$) satisfies $\Delta x\Delta p\geq\nicefrac{{\hbar}}{{2}}$ [17, 18]. This implies that the dimension of a phase-space feature is constrained by this standard limit; if a phase-space feature has dimensions below this threshold, it is considered nonphysical. This general assumption was challenged by Zurek [13], who demonstrated that the sub-Planck structures in the compass state are significant; that is, these spotty features have been found to be crucial in enhancing sensitivity to displacements, and this improvement in sensitivity is directly dedicated to the sub-Planck scale structures in the state. In this section we primarily focus on reviewing the key concept of sub-Planck structures and their critical role in enhancing sensitivity to displacements.
@@ II.1 Phase space and sensitivity
A Schrödinger coherent state can be expressed as a displaced vacuum state |$\ket{\alpha}:=\hat{D}(\alpha)\ket{0}$, where $\hat{D}(\alpha):=\exp(\alpha\hat{a}^{\dagger}-\alpha^{*}\hat{a})$ is the displacement operator [85], with $\alpha\in\mathbb{C}$ and $\hat{a}$ ($\hat{a}^{\dagger}$) the annihilation (creation) operator. Coherent states have properties that mirror classical states [16], but their superposition may exhibit nonclassical features due to the interference phenomena [86]. The Wigner function denoted by $W_{\hat{\rho}}(\bm{\beta})$, with $\bm{\beta}:=(x,p)^{\top}$, constitutes the phase space of a quantum state $\hat{\rho}$ [7], where $x$ and $p$ are the position and momentum pairs, respectively.
The Wigner negativity is one of the primary tools for analyzing the nonclassical characteristics of a state [7, 87, 88, 89, 90]. Mathematically, the Wigner function can be expressed as [91]
@@
The Schrödinger cat state is the superposition of two distinguishable coherent states, and one of the simplest examples of such states is the even cat state [92], which is defined as
|$\ket{\psi}:=\frac{1}{(2+2\text{e}^{-2|\alpha|^{2}})^{\nicefrac{{1}}{{2}}}}\left(\ket{\alpha}+\ket{-\alpha}\right).$ (2)
A cat state is obtained when the parameter $\alpha$ is high enough to maintain the orthogonality criterion $\braket{\alpha}{-\alpha}\approx 0$. If the constituent coherent states in this superposition are not distinguishable, the resulting states are known as Schrödinger kitten states [75]. This generic form of the cat state with the choices ±$\pm\alpha=\nicefrac{{\pm c_{0}}}{{\sqrt{2}}}$ ($c_{0}\in\mathbb{R}$) represents a horizontal cat state aligned along the position axis in the phase space.
We now discuss the concept of sensitivity to phase-space displacement. For a pure quantum state |$\ket{\psi}$, this sensitivity can be mathematically determined by evaluating the overlap function between a quantum state and its slightly displaced version [13, 14, 15, 28]. This involves calculating how much one state resembles another when displaced in phase space, which provides insight into how precisely the quantum state can detect or respond to changes in its phase-space configuration. Mathematically, this sensitivity can be determined by using [14]
@@
with |$\ket{\psi^{\prime}}:=\hat{D}(\delta)\ket{\psi}$. If $S_{\ket{\psi}}(\delta)=0$, then a state and its displaced counterpart are orthogonal for the displacement $\delta$. Here $S_{\ket{\psi}}(\delta)$ is the overlap, with $\delta:=(\delta x,\delta p)^{\top}$, where $\delta x$ and $\delta p$ are values of the displacements applied in the phase space along position ($x$) and momentum ($p$) axis, respectively. The infinitesimal perturbation $\delta$, which makes the perturbed state quasiorthogonal to the initial state, provides information on the sensitivity to displacements. Smaller values of $\delta$ indicate greater sensitivity to displacements.
Figure 2: Central interference of the compass state for (a) $c_{0}=5$ and (b) $c_{0}=8$.
@@
which exhibits a Gaussian form; hence, coherent states are types of Gaussian quantum states [16], as also evident by the corresponding Wigner function of the coherent state shown in Fig. 12(a). Moreover, the phase-space structure of a coherent state follows the minimal limit set by the uncertainty principle, often referred to as the Planck action in phase space. This implies that the phase-space size of a coherent state sets the minimal norm, and the sub-Planck structure is below this limit and can be limited as much as desired by varying the controlling parameter, whereas all of these characteristics are missing in the coherent [47] (the corresponding discussion about the Planck scale is provided in the Appendix A). The sensitivity of a coherent state to displacements in phase space is obtained as
$S_{\ket{\alpha}}(\delta):=\text{e}^{-|\delta|^{2}},$ (5)
@@
which tends to zero for the displacement $|\delta|>1$, implying that the sensitivity of a coherent state falls at the standard limits. The phase-space structure of a coherent state and its sensitivity adhere precisely to the standard quantum-mechanical limits, that is, the coherent state achieves the theoretical minimum uncertainty allowed by the Heisenberg uncertainty principle, reflecting the optimal balance between precision in position and momentum measurements. Consequently, in our analysis, we evaluate each example by comparing it against these established norms. This involves assessing how each example measures up to the theoretical benchmarks and standard limits, allowing us to understand their relative performance and behavior in relation to these reference points.
II.2 Direction-oriented sensitivity enhancement
The Wigner function of a macroscopic horizontal cat state typically manifests as two distinct Gaussian peaks in phase space with a central oscillatory interference pattern directed along the momentum axis, where each Gaussian peak corresponds to a coherent state [93, 94]. Cat states do not exhibit sub-Planck features as their interference phase-space features are not limited in all directions of phase space [47]. Figures 12(b) and 12(c) shown in Appendix A illustrate the central phase-space parts of a horizontal cat state with smaller (kitten state) and larger (macroscopic cat) amplitudes $c_{0}$, respectively. Mathematically, the Wigner function of the horizontal cat state is defined as
$W_{\ket{\psi}}(\beta):=\frac{1}{[2+2\text{exp}(-c_{0}^{2})]}\sum^{2}_{i,j=1}W_{\ket{\alpha_{i}}\bra{\alpha_{j}}}(\beta),$ (6)
where
$-\alpha_{i}\beta^{*}\big{)}\big{]},$ (7)
with
$G_{k,l}:=\exp\bigg{[}-\frac{1}{2}\left(|k|^{2}+|l|^{2}\right)\bigg{]},$ (8)
setting $\alpha_{1}=c_{0}$ and $\alpha_{2}=-c_{0}$ in Eq. (6), with $c_{0}=1$ representing the horizontal kitten state and $c_{0}=5$ representing its larger version. Figures 12(b) and 12(c) in the Appendix A clearly illustrate the difference in phase-space features between the smaller and larger cats, respectively. In comparison to its kitten counterpart and the coherent state, the larger cat state appears to have more fine-grained phase-space features. For example, the central phase-space feature, shown in Fig. 12(c), is constrained along the $p$ direction in the phase space, but it has the same extension along the $x$ direction as those of coherent states, which is highlighted by plotting corresponding zeros indicated with black curves around the patches. This suggests a directional dependence on the finer resolution of the corresponding interference features, which also affects the sensitivity to displacements in the same way. The sensitivity to displacement for a cat state is obtained as
$S_{\ket{\psi}}(\delta):=\left|\sum^{2}_{i,j=1}O_{\ket{\alpha_{i}}\bra{\alpha_{j}}}(\delta)\right|^{2},$ (9)
where
The overlap for $c_{0}=5$ is shown in Fig. 12(i), and it can be verified that the overlap vanishes for $|\delta|<1$ along the momentum axis. Along the position axis in phase space, however, $|\delta|>1$ is needed to make the overlap zero, which aligns with that of a coherent state in the same direction, suggesting that for a cat state, the sensitivity to displacement is heightened along specific phase-space directions. In this case, it is confined along the momentum direction, as shown by the white curves representing the zeros of the overlap function. Additionally, it can be shown that the central structure of the overlap function for the horizontal cat state may contract further along the momentum axis as $c_{0}$ increases, indicating that the sensitivity to displacement improves as $c_{0}$ grows.
Another illustration is the Wigner phase-space features of squeezed-vacuum states, which are limited along the direction of squeezing, implying that in these directions the sensitivity to displacement can be less than that of a coherent state. This shows that squeezed-vaccum states may provide enhanced sensitivity compared to standard coherent states, making them useful in quantum metrological applications [95, 32].
II.3 Sub-Planckian sensitivity